# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v4

@@ 4.4 Ablation studies
We examine the effect of our three key designs (frequency-domain motion decomposition, the alignment loss $\mathcal{L}_{\text{align}}$, and the explicit synchronization) separately. The results after removing each of these three components individually are shown as “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”. Two additional ablations are to remove the two synchronization mechanisms and all three designs, denoted as “w/o $\mathcal{L}_{\text{align}}$, exp sync” and “w/o all”, respectively. More details can be found in Section  C.4. Results in Table 1, 3, 5 show that removing any of the three components can lead to varying extents of performance decline.
Removal of Decomposition. As shown in Figure 10, after removing the decomposition mechanism, once periodic relative motion is involved in the interactions between objects, it becomes easier for high-frequency motions to be neglected, which is intuitively shown as two objects being in an almost relative stationary state, making it difficult to complete the action. This also results in poor performance in semantics metrics FID and RA from Table 1,  3, and  5. In Section  B.3, we’ve provided more detailed experiments to examine the necessity of decomposition in modeling high-frequency components with semantics, through which we have determined that the performance of frequency decomposition significantly surpasses that of merely removing too high-frequency noise.
@@
Removal of $\mathcal{L}_{\text{align}}$ or Explicit Synchronization. As is shown in Figure 7 and 6, removing either one of $\mathcal{L}_{\text{align}}$ or explicit synchronization leads to unreasonable penetration or contact loss between objects or between humans and objects, with explicit synchronization playing a more significant role. This phenomenon is also revealed in quantitative evaluation results from Table  1,  3 and  5. An observed phenomenon is that simply incorporating frequency decomposition (w/o $\mathcal{L}_{\text{align}}$, exp sync) poses higher demand on synchronization, which is manifested as worse contact-based metrics (Defined in Section  4.2) than “w/o all”. Simply integrating $\mathcal{L}_{\text{align}}$ (w/o exp sync) can not fully solve this issue, making explicit synchronization steps become a must.
@@ 4.5 User Study
There are $10$ different splits in our experiments (TACO splits 1-4, CORE4D splits 1-2, OAKINK2 test split, GRAB splits of unseen subjects/objects, BEHAVE test split). We randomly draw 15 samples per split, resulting in a total of $150$ different questions ($15\times 10$). Each question requires participants to choose the best option among Ours and all other baselines, based primarily on “Completion, Realism, and Naturalness”. If the decision is difficult to make, participants are also instructed to consider “additional misalignments such as interpenetration, contact loss, jitter”. $150$ different individuals are involved, and each is assigned $10$ different questions, where the option orders are randomly shuffled. Our mechanism guarantees that each question is answered by exactly $10$ different agents. Figure  9 shows the times each method voted as the “best” in each split.
@@ 5 Conclusions and Discussions
This paper presents SyncDiff, a unified framework for synchronized motion synthesis of multi-body HOI interaction by estimating both data sample scores and alignment scores, and jointly optimizing sample and alignment likelihoods in inference. We also introduce a frequency-domain decomposition to better capture high-frequency motions with semantics. Experiments on five datasets demonstrate that SyncDiff can be adapted to multiple scenarios with any number of humans, hands, and rigid objects. Comparative experiments also demonstrate that in each specific setting, our method achieves better contact accuracy and action semantics than a range of state-of-the-art baselines. Limitations, potential solutions, extensions, and some discussions of SyncDiff are provided in Section  E.
@@ A.1 Diffusion Model Basics
where $10<\beta_{t}<$ 1. We can derive that
@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
1) Illustrate how to define a set of alignment scores featuring synchronization analogous to the commonly used data sample scores in diffusion models, and derive the corresponding loss term $\mathcal{L}_{\text{align}}$ in Section  3.5;
2) Prove that the explicit synchronization formulas in Section  3.6 are equivalent to maximum total likelihood sampling on the newly computed Gaussian distribution, where data sample scores and alignment scores are jointly considered.
Before we start, let’s derive some commonly used formulas in diffusion models that we will need in our proof, which the readers might not be familiar with. If you are familiar with the step-by-step denoising formula of diffusion, you can skip directly to Eq  12 and  13.
@@
A problem here is that $\hat{\mu}$ is predicted by the model based on the result of step $t$, but $(a_{1},b_{1})$, $\dots$, $(a_{K},b_{K})$ all belong to step $(t-1)$ along with $\hat{x}^{{}^{\prime\prime}}$. Here, we make an assumption that $\hat{x}_{t}$ and $\hat{x}_{t-1}$ are close, so that we can take $(a_{1},b_{1})$, $(a_{2},b_{2})$, $\dots$, $(a_{K},b_{K})$ from $\hat{x}_{t}=\hat{x}$.
@@
where $1\leq t_{1}<t_{2}<\dots<t_{R}\leq T$ are the synchronization timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent standard variances(without synchronization). The value of hyperparameter $\lambda_{\text{exp}}$ can be found in Table  17. Therefore,
@@ B.1 Determine the cutoff boundary $L$ for Frequency Decomposition
In frequency decomposition (Section  3.3), we discard signals with frequencies higher than $\phi_{L}/2\pi$, where $\phi_{L}=L/N$, and $L$ is the cutoff boundary. We analyzed the impact of different $L$ on the maximum error $\epsilon$, which is averaged on the dimension of $N$ (number of frames) and takes the maximum across all degrees of freedom (the dimension of $D_{\text{sum}}$). The errors listed in Table  6 are averaged over all data samples in the datasets and are measured in millimeters.
@@
Since $L=16$ is already enough to almost perfectly reconstruct the motion sequence, to strike a balance between motion representation fidelity and filtering of too high-frequency noise in mocap datasets, in practice, we take $L=16$.
B.2 Determine the Interval $s$ for Explicit Synchronization
In explicit synchronization (Section  3.6), to balance efficiency and performance, we choose a hyperparameter $s(s<<T)$, where $T=1000$ is the total number of denoising steps, which means we only perform explicit synchronization operations every $s$ step, $R=T/s$ times in total.
@@ B.3 Check the Reconstruction Quality of High-frequency Components
Identifying the source of less jitter is relatively straightforward. It is noted that the motion trajectories for training in the “w/o decompose” scenario includes those too high-frequency noise (See Section  C.4). However, as observed in the “w/o decompose” videos of TACO, there is almost no large-scale jitter in the synthesized motions, with the only issue being the absence of high-frequency interactions between objects. This is because, in the “w/o decompose” experiments, both synchronization mechanisms are effective, ensuring that the generated motions are sufficiently aligned and thus reducing jitter. Conversely, in the OAKINK2 ablation studies of synchronization, whether it’s “w/o align loss” or “w/o exp sync,” the use of frequency decomposition also removes too high-frequency noise, yet it still results in greater jitter compared to the complete SyncDiff (Ours). This sufficiently demonstrates that the contribution to less jitter is primarily due to the synchronization mechanisms, rather than the filter of too high-frequency noise. A more straightforward method of verification is to observe that the errors in Table  6 caused by the filter of too high-frequency noise are almost negligible compared to the scales of the motions, indicating that this operation theoretically has minimal impact on the model’s performance.
@@
Experiment 1: On the four test splits of TACO, we only add noise for no larger than $200$ steps (where the full train pipeline needs $1000$ steps), and examine the reconstruction ability of different methods. For any predicted motion trajectory $\hat{x}$, we decompose it as described in Section  3.3, to obtain the high-frequency component $\hat{x}_{\text{ac}}$. We then examine the discrepancy between $\hat{x}_{\text{ac}}$ and the ground truth $x_{\text{ac}}$. The experimental results on the four splits of TACO are shown in Table  8, proving that explicitly representing high-frequency signals with semantics in the frequency domain indeed improves the effect of reconstruction.
@@
Experiment 2: We test the semantic quality of synthesized motions on the four test splits of TACO. The experimental setting is the same as Table  1 in the main text. Results are shown in Table  9. We can clearly see that frequency decomposition enhances the semantic quality of the synthesized motion, and this can not be achieved by merely removing too high-frequency noise.
@@ C.1 Formulas for rel and comb
In Section  3.5,  3.6, for two distinct bodies a and b, we define $\text{rel}(x_{\textbf{a}},x_{\textbf{b}})$ as b’s motion relative to a, and let $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilize the individual motion of a and relative motion between b and a to compute b’s motion. The detailed expressions are shown below:
If $\textbf{a}=o_{j_{1}}$ and $\textbf{b}=o_{j_{2}}$ are two rigid objects, whose motions are $x_{o_{j_{1}}}=[\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}]$ and $x_{o_{j_{2}}}=[\textbf{t}_{o_{j_{2}}},\textbf{q}_{o_{j_{2}}}]$, denote $x_{o_{j_{2}}\rightarrow o_{j_{1}}}=[\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}},\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}]$, then
$\text{rel}(x_{o_{j_{1}}},x_{o_{j_{2}}})=\left[\textbf{q}_{o_{j_{1}}}^{-1}(\textbf{t}_{o_{j_{2}}}-\textbf{t}_{o_{j_{1}}}),\textbf{q}_{o_{j_{1}}}^{-1}\textbf{q}_{o_{j_{2}}}\right],$ (18)
$\text{comb}(x_{o_{j_{1}}},x_{o_{j_{2}}\rightarrow o_{j_{1}}})=[\textbf{q}_{o_{j_{1}}}\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}}+\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}].$ (19)
If $\textbf{a}=o_{j}$ is rigid object, and $\textbf{b}=h_{i}$ is an articulated skeleton, let $x_{o_{j}}=[\textbf{t}_{j},\textbf{q}_{j}]$, and denote the motion for one of the joints in $x_{h_{i}}$ as $\textbf{p}_{h_{i}}$, the motion for one of the joints in $x_{h_{i}\rightarrow o_{j}}$ as $\textbf{p}_{h_{i}\rightarrow o_{j}}$. We have
$\text{rel}(x_{o_{j}},\textbf{p}_{h_{i}})=\textbf{q}_{j_{1}}^{-1}(\textbf{p}_{h_{i}}-\textbf{t}_{j_{1}}),$ (20)
$\text{comb}(x_{o_{j}},\textbf{p}_{h_{i}\rightarrow o_{j}})=\textbf{q}_{o_{j}}\textbf{p}_{h_{i}\rightarrow o_{j}}+\textbf{t}_{o_{j}}.$ (21)
Note that a must be a rigid object, in order to define a coordinate system based on its transformation matrix. More details can be found in the second question of Section  E.1. Also, $\text{rel}\left(\hat{x}_{\textbf{a}},\hat{x}_{\textbf{b}}\right)$ is different from $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$. The former is calculated based on the predicted motion of a and b, which should approach $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$ in order to foster synchronization. The latter is directly predicted by the diffusion model, which should fit the distribution of ground-truth $x_{\textbf{b}\rightarrow\textbf{a}}$, in order to increase data fidelity. Similar arguments hold for comb.
@@ C.2 Post Process: Mesh Reconstruction
SMPL-X Human Body Mesh Reconstruction. Similar to MANO in hand pose representation, for human body pose representation, we also have SMPL-X [57] representation. It is composed of a set of parameters $(\theta,R,l,\theta_{\text{left}},\theta_{\text{right}})$, where the body pose $\theta\in\mathbb{R}^{N\times 21\times 3}$ represents the $21$ joint orientations(except root), $R,l\in\mathbb{R}^{N\times 3}$ are global orientation and translation, and $\theta_{\text{left}},\theta_{\text{right}}\in\mathbb{R}^{N\times 12}$ are compressed representations of two hands. Human shape parameters $\beta\in\mathbb{R}^{10}$ are given, while $\hat{x}\in\mathbb{R}^{N\times 22\times 3}$ denotes the joint positions we predict.
@@
The hyperparameters involved in the mesh reconstruction process can be found in Table  10 and  11.
@@
Optimizer AdamW, $\text{lr}=0.01$
Epoch $5$k $8$k $5$k
$\lambda_{\text{pos}}$ $1$ $1$ $1$
$\lambda_{\text{angle}}$ $0.2$ $0.2$ $0.05$
$\lambda_{\text{vel}}$ $0.03$ $0.03$ $0.02$
@@
Optimizer AdamW, $\text{lr}=0.001$
Epoch $5$k $2$k
$\lambda_{\text{pos}}$ $3$ $1$
$\lambda_{\text{vel}}$ $0.1$ $0.1$

### v5

@@ 4.4 Ablation studies
We examine the effect of our three key designs (frequency-domain motion decomposition, the alignment loss $\mathcal{L}_{\text{align}}$, and the explicit synchronization) separately. The results after removing each of these three components individually are shown as “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”. Two additional ablations are to remove the two synchronization mechanisms and all three designs, denoted as “w/o $\mathcal{L}_{\text{align}}$, exp sync” and “w/o all”, respectively. More details can be found in Section C.4. Results in Tables 1, 3, 5 show that removing any of the three components can lead to varying extents of performance decline.
Removal of Decomposition. As shown in Figure 10, after removing the decomposition mechanism, once periodic relative motion is involved in the interactions between objects, it becomes easier for high-frequency motions to be neglected, which is intuitively shown as two objects being in an almost relative stationary state, making it difficult to complete the action. This also results in poor performance in semantics metrics FID and RA from Tables 13, and 5. In Section B.3, we’ve provided more detailed experiments to examine the necessity of decomposition in modeling high-frequency components with semantics, through which we have determined that the performance of frequency decomposition significantly surpasses that of merely removing too high-frequency noise.
@@
Removal of $\mathcal{L}_{\text{align}}$ or Explicit Synchronization. As is shown in Figures 6 and 7, removing either one of $\mathcal{L}_{\text{align}}$ or explicit synchronization leads to unreasonable penetration or contact loss between objects or between humans and objects, with explicit synchronization playing a more significant role. This phenomenon is also revealed in quantitative evaluation results from Tables 13 and 5. An observed phenomenon is that simply incorporating frequency decomposition (w/o $\mathcal{L}_{\text{align}}$, exp sync) poses higher demand on synchronization, which is manifested as worse contact-based metrics (Defined in Section 4.2) than “w/o all”. Simply integrating $\mathcal{L}_{\text{align}}$ (w/o exp sync) can not fully solve this issue, making explicit synchronization steps become a must.
@@ 4.5 User Study
There are 101010 different splits in our experiments (TACO splits 1-4, CORE4D splits 1-2, OAKINK2 test split, GRAB splits of unseen subjects/objects, BEHAVE test split). We randomly draw 15 samples per split, resulting in a total of 150150150 different questions (15×1015$\times$ 1015 × 10). Each question requires participants to choose the best option among Ours and all other baselines, based primarily on “Completion, Realism, and Naturalness”. If the decision is difficult to make, participants are also instructed to consider “additional misalignments such as interpenetration, contact loss, jitter”. 150150150 different individuals are involved, and each is assigned 101010 different questions, where the option orders are randomly shuffled. Our mechanism guarantees that each question is answered by exactly 101010 different agents. Figure 9 shows the times each method voted as the “best” in each split.
@@ 5 Conclusions and Discussions
This paper presents SyncDiff, a unified framework for synchronized motion synthesis of multi-body HOI interaction by estimating both data sample scores and alignment scores, and jointly optimizing sample and alignment likelihoods in inference. We also introduce a frequency-domain decomposition to better capture high-frequency motions with semantics. Experiments on five datasets demonstrate that SyncDiff can be adapted to multiple scenarios with any number of humans, hands, and rigid objects. Comparative experiments also demonstrate that in each specific setting, our method achieves better contact accuracy and action semantics than a range of state-of-the-art baselines. Limitations, potential solutions, extensions, and some discussions of SyncDiff are provided in Section E.2.
@@ A.1 Diffusion Model Basics
where $0<\beta_{t}<1$. We can derive that
@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
1) Illustrate how to define a set of alignment scores featuring synchronization analogous to the commonly used data sample scores in diffusion models, and derive the corresponding loss term $\mathcal{L}_{\text{align}}$ in Section 3.5;
2) Prove that the explicit synchronization formulas in Section 3.5 are equivalent to maximum total likelihood sampling on the newly computed Gaussian distribution, where data sample scores and alignment scores are jointly considered.
Before we start, let’s derive some commonly used formulas in diffusion models that we will need in our proof, which the readers might not be familiar with. If you are familiar with the step-by-step denoising formula of diffusion, you can skip directly to Eq  9 and  10.
@@
[…]
@@
where $1\leq t_{1}<t_{2}<\dots<t_{R}\leq T$ are the synchronization timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent standard variances(without synchronization). The value of hyperparameter $\lambda_{\text{exp}}$ can be found in Table  19. Therefore,
@@ B.1 Determine the cutoff boundary $L$ for Frequency Decomposition
In frequency decomposition of Section 3.3, we discard signals with frequencies higher than $\phi_{L}/2\pi$, where $\phi_{L}=L/N$, and $L$ is the cutoff boundary. We analyzed the impact of different $L$ on the maximum error $\epsilon$, which is averaged on the dimension of $N$ (number of frames) and takes the maximum across all degrees of freedom (the dimension of $D_{\text{sum}}$). The errors listed in Table  6 are averaged over all data samples in the datasets and are measured in millimeters.
@@
Our goal is to minimize $L$, provided that the error between filtered and raw motions remains negligible relative to the original signal amplitude. To strike a balance between motion representation fidelity and simplicity, as well as filtering of too high-frequency noise in mocap datasets, in practice, we take $L=16$.
B.2 Determine the Interval $s$ for Explicit Synchronization and Computational Cost for the Process
In explicit synchronization from Section 3.6, to balance inference speed and performance, we choose a hyperparameter $s(s<<T)$, where $T=1000$ is the total number of denoising steps, which means we only perform explicit synchronization operations every $s$ step, $R=T/s$ times in total.
@@ B.2 Determine the Interval $s$ for Explicit Synchronization and Computational Cost for the Process
Here we also list the computational cost comparison between SyncDiff and each baselines. We test inference speed on a single NVIDIA A40 GPU for synthesizing a 200-frame sequence. Time consumption across different methods are shown in Tables  8 and  9.
Method Dataset CORE4D BEHAVE
OMOMO 7.1s 5.9s
CG-HOI 10.7s 7.6s
SyncDiff (Ours) 6.5s 4.2s
w/o exp sync 5.4s 3.6s
Table 8: Inference time for a 200-frame human-object-interaction sequence.
Method Dataset TACO OAKINK2 GRAB
MACS 7.4s 7.0s 5.7s
DiffH2O 8.6s 8.9s 7.4s
SyncDiff (Ours) 7.7s 7.3s 6.6s
w/o exp sync 5.9s 5.6s 5.4s
Table 9: inference time for a 200-frame hand-object-interaction sequence.
[…]
@@ B.3 Check the Reconstruction Quality of High-frequency Components
Identifying the source of less jitter is relatively straightforward. It is noted that the motion trajectories for training in the “w/o decompose” scenario includes those too high-frequency noise (See Section C.4). However, as observed in the “w/o decompose” videos of TACO, there is almost no large-scale jitter in the synthesized motions, with the only issue being the absence of high-frequency interactions between objects. This is because, in the “w/o decompose” experiments, both synchronization mechanisms are effective, ensuring that the generated motions are sufficiently aligned and thus reducing jitter. Conversely, in the OAKINK2 ablation studies of synchronization, whether it’s “w/o align loss” or “w/o exp sync,” the use of frequency decomposition also removes too high-frequency noise, yet it still results in greater jitter compared to the complete SyncDiff (Ours). This sufficiently demonstrates that the contribution to less jitter is primarily due to the synchronization mechanisms, rather than the filter of too high-frequency noise. A more straightforward method of verification is to observe that the errors in Table  6 caused by the filter of too high-frequency noise are almost negligible compared to the scales of the motions, indicating that this operation theoretically has minimal impact on the model’s performance.
@@
Experiment 1: On the four test splits of TACO, we only add noise for no larger than 200200200 steps (where the full train pipeline needs 100010001000 steps), and examine the reconstruction ability of different methods. For any predicted motion trajectory $\hat{x}$, we decompose it as described in Section 3.3, to obtain the high-frequency component $\hat{x}_{\text{ac}}$. We then examine the discrepancy between $\hat{x}_{\text{ac}}$ and the ground truth $x_{\text{ac}}$. The experimental results on the four splits of TACO are shown in Table  10, proving that explicitly representing high-frequency signals with semantics in the frequency domain indeed improves the effect of reconstruction.
@@
Experiment 2: We test the semantic quality of synthesized motions on the four test splits of TACO. The experimental setting is the same as Table 1 in the main text. Results are shown in Table  11. We can clearly see that frequency decomposition enhances the semantic quality of the synthesized motion, and this can not be achieved by merely removing too high-frequency noise.
@@ C.1 Formulas for rel and comb
In Sections 3.53.6, for two distinct bodies a and b, we define $\text{rel}(x_{\textbf{a}},x_{\textbf{b}})$ as b’s motion relative to a, and let $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilize the individual motion of a and relative motion between b and a to compute b’s motion. The detailed expressions are shown below:
If $\textbf{a}=o_{j_{1}}$ and $\textbf{b}=o_{j_{2}}$ are two rigid objects, whose motions are $x_{o_{j_{1}}}=[\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}]$ and $x_{o_{j_{2}}}=[\textbf{t}_{o_{j_{2}}},\textbf{q}_{o_{j_{2}}}]$, denote $x_{o_{j_{2}}\rightarrow o_{j_{1}}}=[\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}},\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}]$, then $\text{rel}(x_{o_{j_{1}}},x_{o_{j_{2}}})=\left[\textbf{q}_{o_{j_{1}}}^{-1}(\textbf{t}_{o_{j_{2}}}-\textbf{t}_{o_{j_{1}}}),\textbf{q}_{o_{j_{1}}}^{-1}\textbf{q}_{o_{j_{2}}}\right]$, $\text{comb}(x_{o_{j_{1}}},x_{o_{j_{2}}\rightarrow o_{j_{1}}})=[\textbf{q}_{o_{j_{1}}}\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}}+\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}]$.
If $\textbf{a}=o_{j}$ is rigid object, and $\textbf{b}=h_{i}$ is an articulated skeleton, let $x_{o_{j}}=[\textbf{t}_{j},\textbf{q}_{j}]$, and denote the motion for one of the joints in $x_{h_{i}}$ as $\textbf{p}_{h_{i}}$, the motion for one of the joints in $x_{h_{i}\rightarrow o_{j}}$ as $\textbf{p}_{h_{i}\rightarrow o_{j}}$. We have $\text{rel}(x_{o_{j}},\textbf{p}_{h_{i}})=\textbf{q}_{j_{1}}^{-1}(\textbf{p}_{h_{i}}-\textbf{t}_{j_{1}})$, $\text{comb}(x_{o_{j}},\textbf{p}_{h_{i}\rightarrow o_{j}})=\textbf{q}_{o_{j}}\textbf{p}_{h_{i}\rightarrow o_{j}}+\textbf{t}_{o_{j}}$.
Note that a must be a rigid object, in order to define a coordinate system based on its transformation matrix. Also, $\text{rel}\left(\hat{x}_{\textbf{a}},\hat{x}_{\textbf{b}}\right)$ is different from $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$. The former is calculated based on the predicted motion of a and b, which should approach $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$ in order to foster synchronization. The latter is directly predicted by the diffusion model, which should fit the distribution of ground-truth $x_{\textbf{b}\rightarrow\textbf{a}}$, in order to increase data fidelity. Similar arguments hold for comb.
@@ C.2 Post Process: Mesh Reconstruction
SMPL-X Human Body Mesh Reconstruction. Similar to MANO in hand pose representation, for human body pose representation, we also have SMPL-X [57] representation. It is composed of a set of parameters $(\theta,R,l,\theta_{\text{left}},\theta_{\text{right}})$, where the body pose $\theta\in\mathbb{R}^{N\times 21\times 3}$ represents the 212121 joint orientations(except root), $R,l\in\mathbb{R}^{N\times 3}$ are global orientation and translation, and $\theta_{\text{left}},\theta_{\text{right}}\in\mathbb{R}^{N\times 12}$ are compressed representations of two hands. Human shape parameters $\beta\in\mathbb{R}^{10}$ are given, while $\hat{x}\in\mathbb{R}^{N\times 22\times 3}$ denotes the joint positions we predict.
@@
The hyperparameters involved in the mesh reconstruction process can be found in Table  12 and  13.
@@
[…]
Epoch 555k 888k 555k
$\lambda_{\text{pos}}$ 111 111 111
$\lambda_{\text{angle}}$ 0.20.20.2 0.20.20.2 0.050.050.05
$\lambda_{\text{vel}}$ 0.030.030.03 0.030.030.03 0.020.020.02
@@
[…]
Epoch 555k 222k
$\lambda_{\text{pos}}$ 333 111
$\lambda_{\text{vel}}$ 0.10.10.1 0.10.10.1